# Developers’ Perspectives on Software Licensing: Current Practices, Challenges, and Tools

### v1

@@ Abstract.
Most modern software products incorporate open-source components, requiring development teams to maintain compliance with each component’s licenses. Noncompliance can lead to significant financial, legal, and reputational repercussions. While some organizations may seek advice from legal practitioners to assist with licensing tasks, developers still play a key role in such a process. To this end, it is essential to understand how developers approach license compliance tasks, the challenges they encounter, and the tools that they use. This work studies these aspects of software licensing practices through a study—conducted by a joint team of software engineering and legal researchers—consisting of a survey with 58 software developers and seven follow-up interviews. The study resulted in ???? key findings regarding the current state of practice. We discuss the implications of our findings and offer directions for future research as well as actionable recommendations for licensing tools.
@@ 1. Introduction
Software reuse, particularly through open source software (OSS), has become fundamental in the process of software development, allowing software developers to avoid duplicating efforts in areas where previous developers have found suitable solutions. However, the fact that a component is open source does not mean that it can be used without restrictions. Because software is protected by copyright and patent law (cop, 2021), OSS components are typically released under one or more OSS licenses that govern how a component may be reused, modified, and distributed. Failure to comply with the terms of any applicable license risks liability for infringement, which can result in legal, financial, and reputational consequences (Claburn, 2022; Vincent, 2022; Claburn, 2023; Gunningham et al., 2004; Vance, 2010; Gross, 2007), both for individual developers and for the organizations for which they work.
@@
Ensuring license compliance throughout the development process is, therefore, a critical task. However, this process can be challenging to manage. Software systems can involve hundreds or thousands of OSS components, each under a separate license that must be identified and understood (Ballhausen, 2019). Furthermore, some OSS licenses have conflicting terms, making the use of components under different licenses problematic. There are hundreds of potentially applicable OSS licenses (spd, 2024; osi, 2024), some of which use legal terminology that is unfamiliar to developers and that does not always have a broadly accepted interpretation (Almeida et al., 2019; Gangadharan et al., 2012; Vendome et al., 2018).
@@
As a result, it may seem appropriate to offload the work of license compliance to legal teams or Open Source Program Offices (OSPOs), but this may delay compliance efforts until near the end of a development cycle (Wintersgill et al., 2024), leading to lower productivity and wasted effort. Thus, while legal experts play an important role in compliance processes, it is also important for developers to be trained in such tasks.
@@
As the individuals closest to the code, developers bear responsibility for maintaining compliance with the licenses governing that code (Wintersgill et al., 2024). However, where previous work has identified that licensing tasks can be challenging (Almeida et al., 2019, 2017; Vendome et al., 2018; Moraes et al., 2021; German and Hassan, 2009; Gangadharan et al., 2012), it does not identify developers’ perceptions of different licensing tasks, and questions remain regarding developers’ adoption of licensing tools and the strategies employed when addressing licensing issues.
To fill this knowledge gap, this study surveyed a variety of software developers, including those who may not perform compliance tasks regularly. Our goal was to gain a deeper understanding of the licensing processes and challenges for developers, discover the gaps in the technology developers use for licensing, and work toward organizational and technical solutions.
@@
  • A novel assessment of how developers not only understand but engage with software licensing tasks, based on a survey of 58 developers with various levels of experience in licensing;
@@ 2.1. Backgrounds on Software licensing
As software is protected by copyright law (among other legal doctrines) in the United States and around the world (cop, 2021), the owner of a piece of software’s copyright holds the rights to distribute, reproduce, or create derivative works of that software, as well as the right to authorize others to do so (USc, 2021). Such authorization often occurs by means of a license, which sets forth the permissible uses of the software and includes any restrictions on use. (Meeker, 2017) Use outside of the scope of the license constitutes copyright infringement, unless such use constitutes a fair use under U.S. copyright law.
@@
OSS developers often choose one of several popular existing licenses to apply to their software, such as the MIT (mit, [n. d.]) license or a member of the GPL (gpl, [n. d.]) family of licenses. Permissive licenses, as the name suggests, typically allow broad use of the software, requiring only that, for example, the user provide notice files with information on the license or on attribution. Copyleft licenses, by contrast, constrain future use of the software by, typically, requiring that the user make the full source code of their work available under the same license. A developer must therefore understand, at a minimum, which licenses apply to components they incorporate into their work and what those licenses require.
@@
Although several licenses have been widely used for many years, interpretive questions still exist. Besides the preliminary questions about the types of uses a license permits or prohibits, questions can arise about whether two licenses are compatible with each other, as well as how certain terms apply to desired software uses. For example, the requirements of GPL v2 apply to any work “based on the Program,” which the license defines as equivalent to “any derivative work under copyright law.” But this language does not answer the question of whether system calls to a library create a work “based on the Program” or whether the answer depends on whether the link is static or dynamic, a question that even OSS lawyers have identified as unresolved  (Wintersgill et al., 2024).
2.2. Developers and licensing
@@
Previous work has shown that software developers often find licensing tasks difficult, such as determining whether two licenses are compatible with each other (Almeida et al., 2019, 2017), ensuring that their software complies with all relevant licensing requirements (Vendome et al., 2018), and releasing their software under multiple licenses (Moraes et al., 2021). Building on this work, we seek to contextualize these challenges by examining how often they are encountered and how developers address them.
@@
Previous work also shows us that several methods exist to assist with licensing tasks. Both open source and proprietary tools can help with license compliance, such as Fossology (Fos, [n. d.]) and Black Duck (Bla, [n. d.]), respectively. Software bills of materials (SBOMs) can provide insights into software’s composition, facilitating licensing tasks (Bi et al., 2024; Stalnaker et al., 2024). Additionally, researchers have developed a variety of tools and processes aimed at assisting developers with compliance tasks (Tuunanen et al., 2009; Ombredanne, 2020; Kapitsaki et al., 2017; German and Hassan, 2009; German and Di Penta, 2012; German et al., 2010; Hemel et al., 2011; Feng et al., 2019; Gangadharan et al., 2008). Some of these studies have explored the use of generative AI to assist in compliance tasks (Cui et al., 2025; Kahol et al., 2025; Ke et al., 2025); Other studies, based on mined software repositories and Q&A websites, have explored methods to detect non-approved licenses, non-standard license variants, and exceptions.  (Meloca et al., 2018; Zacchiroli, 2022; Papoutsoglou et al., 2022; Vendome et al., 2017b; Xu et al., 2023). Where previous work has sought to uncover ways to address licensing issues, our work seeks to understand if and how developers are utilizing these methods.
@@
Prior work has also explored some of the processes and decisions developers make during license compliance efforts. Wu et al. conducted a large empirical study aimed at understanding license usage (Wu et al., 2024). Works by Vendome et al. and Di Penta et al. sought to understand when, how, and why developers opt to change the license of their software (Di Penta et al., 2010; Vendome et al., 2017a). Liu et al. introduced a method to predict license changes from source code changes (Liu et al., 2019). Relatedly, Wintersgill et al. surveyed and interviewed legal professionals with experience in software license compliance to understand their challenges and their views on software license compliance tasks (Wintersgill et al., 2024). Lastly, Li et al. provide a systematic literature review of stakeholders’ challenges in license compliance tasks (Li et al., 2025).
@@
Where the literature offers insights, solutions, and tools for licensing tasks, this work seeks to expand on previous efforts in two key ways: (1) by understanding if and how developers adopt these tools and techniques in practice, and (2) by identifying the license compliance pain points that are most relevant to developers by providing greater insight into the specific problems developers report encountering most often, which they report are most difficult, and what strategies they use to address them.
@@ 3. Study Methodology
The goal of this study is to analyze the processes employed and challenges faced by software developers in performing OSS license compliance tasks. The study was conducted through a survey and follow-up interviews, and involved a collaboration between software engineering (SE) and law researchers. The context of the study consists of 58 developers recruited through personal contacts and open-source projects’ discussion channels.
More specifically, the study aims to address the following three research questions (RQs):
@@
    What are the licensing challenges faced by software developers in practice? This RQ explores the licensing compliance challenges developers face in their routine workflows.
@@
    What are the processes by which software licensing problems are resolved? This RQ complements the previous one by exploring how developers resolve compliance challenges when they occur, and investigating how developers perceive the relationship between development and legal teams.
@@
    How are developers incorporating license compliance tools in their workflows, if at all, and what issues have they encountered? Here we explore how licensing tools are being used by developers, if at all, as well as what issues and shortcomings developers report when using such tools.
This work, including the survey questionnaire, the participant identification procedure, and the survey and interview protocols, was approved by our institution’s ethics review board. An overview of our methodology is depicted in Figure 1.
Figure 1. Research methodology (image credits at (Nathan Wintersgill, Trevor Stalnaker, Daniel Otten, Laura A. Heymann, Oscar Chaparro, Massimiliano Di Penta, Daniel M. German, Denys Poshyvanyk, 2025))
@@ 3.1. Survey design and participant identification
When designing our survey questionnaire, we considered previous literature on software license compliance, as well as general guidelines (Groves et al., 2009) and SE-specific guidelines for survey design (Pfleeger and Kitchenham, 2001; Kitchenham and Pfleeger, 2002a, b, c, d, 2003). The survey went through multiple review cycles, involving seven SE researchers and one law researcher, to achieve appropriate content, clear and concise language, and eliminate confusion and bias. We sought to limit the response time of the survey to minimize the cognitive load on participants and to encourage higher completion rates. In addition, we conducted a small pilot study with members of our research laboratory to further refine the survey questions, including by identifying questions that could be culled or reworked to better focus on the study’s research goals. The final survey included questions to obtain information related to each of our RQs, including questions relating to the compliance challenges faced by developers, existing compliance practices, and the use of tooling or other automation to perform compliance. Figure 2 provides an overview of our survey questionnaire. The complete survey text can be found in our online replication package (Nathan Wintersgill, Trevor Stalnaker, Daniel Otten, Laura A. Heymann, Oscar Chaparro, Massimiliano Di Penta, Daniel M. German, Denys Poshyvanyk, 2025).
@@
To recruit developers with diverse experiences, we distributed the survey in three ways: (1) through our professional networks (e.g., posts on LinkedIn and emails), (2) through snowball sampling, encouraging respondents to share the survey within their networks, and (3) by posting links to the survey on relevant servers on Discord.
To identify Discord communities, we examined the TIOBE index (Jansen, 2025) of top programming languages to identify the top 20 most popular programming languages as of January 2025. We then searched for Discord servers associated with these languages (excluding Scratch, which is primarily a tool for beginners) by locating server links through the languages’ own documentation or websites, through third-party Discord server browsing tools DiscordMe (DiscordMe, [n. d.]) and Disboard (Disboard, [n. d.]), and through Reddit posts on subreddits related to the targeted languages. This process was further supplemented by adding three generalized programming/development-focused servers and one server each for the popular React and Node.js frameworks. Altogether, this yielded 27 Discord servers representing 16 of the top programming languages as well as general programming spaces and popular frameworks. Next, one researcher carefully analyzed the rules of each Discord server to determine whether it would permit a solicitation for participation in a survey, asking server administrators in cases where an answer was not immediately clear. In total, we were able to post survey invitations in 14 software development-focused Discord servers, with total membership exceeding 485k users at the time of distribution. Two additional posts were made in smaller servers but were later removed by moderators.
@@ 3.2. Survey response collection and analysis
Survey responses were collected using Qualtrics (qua, nd) from February 2025 to June 2025. Survey invitations were sent out iteratively during this same period. To monitor where responses originated from, we distributed survey URLs with identifiers unique to each source of participants. We obtained 58 complete survey responses. Of these, 38 came from our professional networks and 17 came from Discord, while 3 participants seemingly removed the tracking component from the survey URL. The largest population from a single Discord server was 6 participants.
Responses to the five open-ended questions were analyzed through a qualitative coding approach (Spencer, 2009). Three authors, all SE researchers, performed open-coding by independently assigning one or more codes to each response using a shared spreadsheet and codebook. Each annotator independently coded all 58 responses, adding new codes to the codebook as necessary. Once the initial coding was completed, the annotators met to settle disagreements and consolidate the set of codes. Reasons for disagreement included different interpretations of potentially ambiguous responses, applying the same codes in different contexts, and using broader codes rather than more specific ones. All such disagreements were resolved among the annotators, resulting in a final set of codes, definitions, and assignments for each response. Our replication package (Nathan Wintersgill, Trevor Stalnaker, Daniel Otten, Laura A. Heymann, Oscar Chaparro, Massimiliano Di Penta, Daniel M. German, Denys Poshyvanyk, 2025) contains our final codes and definitions. We did not base our analysis on inter-rater agreements, as the nature of the study was such that codes were developed through an inductive process, and multiple codes could be assigned to each survey response. However, we carefully followed best open-coding practices (Spencer, 2009), and we leveraged coders’ discussions to ensure the reliability of the results.
We have corrected grammar and spelling errors in quotations from open-ended questions for readability and have indicated by ellipses or brackets any deletions or additions to text for clarity or space-related reasons. All responses have been attributed to survey participants using anonymous IDs for traceability (e.g., $R_{12}$).

### v2

@@ Abstract.
Most modern software products incorporate open-source components, requiring development teams to maintain compliance with each component’s licenses. Noncompliance can have significant financial, legal, and reputational repercussions. Although some organizations may seek advice from legal practitioners to assist with licensing tasks, developers still play a key role in this process. To this end, it is essential to understand how developers approach licensing compliance tasks, the challenges they encounter, and the tools they use. This work studies these aspects of software licensing practices through a study—conducted by a joint team of software engineering and legal researchers—consisting of a survey with 58 software developers and 7 follow-up interviews. The study resulted in ???? key findings regarding the current state of practice. We discuss the implications of our findings and offer directions for future research, as well as actionable recommendations.
@@ 1. Introduction
Software reuse, particularly through open source software (OSS), has become fundamental in the process of software development, allowing software developers to avoid duplicating efforts in areas where previous developers have found suitable solutions. However, the fact that a component is open source does not mean that it can be used without restrictions. Because software is protected by copyright and patent law (8), OSS components are typically released under one or more OSS licenses that govern how a component may be reused, modified, and distributed. Failure to comply with the terms of any applicable license risks liability for infringement, which can result in legal, financial, and reputational consequences (Claburn, 2022; Vincent, 2022; Claburn, 2023; Gunningham et al., 2004; Vance, 2010; Gross, 2007), both for individual developers and for the organizations for which they work.
@@
Ensuring license compliance throughout the development process is therefore a critical task. However, this process can be challenging to manage. Software systems can involve hundreds or thousands of OSS components, each under a separate license that must be identified and understood (Ballhausen, 2019). Furthermore, some OSS licenses have conflicting terms, making the use of components under different licenses problematic. There are hundreds of potentially applicable OSS licenses (48; 44), some of which use legal terminology that is unfamiliar to developers and that does not always have a broadly accepted interpretation (Almeida et al., 2019; Gangadharan et al., 2012; Vendome et al., 2018).
@@
As a result, it may seem appropriate to offload the license compliance work to legal teams or Open Source Program Offices (OSPOs), but this can delay compliance efforts until near the end of a development cycle (Wintersgill et al., 2024), leading to lower productivity and wasted effort. Thus, while legal experts play an important role in compliance processes, it is also important for developers to be trained in such tasks.
@@
As the individuals closest to the code, developers bear primary responsibility for maintaining compliance with the licenses governing that code (Wintersgill et al., 2024). Although previous work has identified that licensing tasks can be challenging for developers (Almeida et al., 2019, 2017; Vendome et al., 2018; Moraes et al., 2021; German and Hassan, 2009; Gangadharan et al., 2012), further investigation of developers’ perceptions of different licensing tasks is needed, and questions remain regarding developers’ use of licensing tools and the strategies employed to address licensing tasks.
To fill this knowledge gap, this study surveyed software developers with various levels of familiarity with licensing tasks. Given that licensing tasks might not need to be performed as frequently as other software engineering tasks and that developers do not always have training regarding software licensing, this study provides important context by representing the views of developers with realistic licensing workloads and training, spanning those with limited exposure to licensing tasks to those who address such tasks frequently, which results in a novel view of the state of practice. Our goal was to gain a deeper understanding of licensing processes and challenges for developers, discover the gaps in the technology developers use for licensing, and work toward organizational and technical solutions. We accomplished this by identifying the licensing tasks that developers perform most frequently and the tasks that are most difficult, explaining how developers address licensing tasks through strategies that encompass both communication with legal experts and making changes themselves, and diagnosing low awareness and adoption of tools that can assist with licensing tasks.
@@
  • A novel assessment of how developers not only understand but also engage with software licensing tasks, based on a survey of 58 developers with various levels of experience in licensing;
@@ 2.1. Software licensing
As software is protected by copyright law (among other legal doctrines) in the United States and around the world (8), the owner of a piece of software’s copyright holds the rights to distribute, reproduce, and create derivative works of that software, as well as the right to authorize others to do so (52). Such authorization often occurs by means of a license, which can grant permissions such as using, modifying, or redistributing the software, as well as potentially applying restrictions to those permissions. (Meeker, 2017) Use outside of the scope of the license constitutes copyright infringement, unless such use constitutes a fair use under U.S. copyright law.
@@
OSS developers often choose one of several existing popular licenses to apply to their software, such as the MIT (39) license or a member of the GPL (21) family of licenses. Permissive licenses, as the name suggests, typically allow broad use of the software, requiring only that, for example, the user provide notice files with information on the license or on attribution. Copyleft licenses, by contrast, constrain future use of the software by, typically, requiring that the user make the full source code of their work available under the same license. A developer must therefore understand, at a minimum, which licenses apply to components they incorporate into their work and what those licenses require if they wish to remain compliant and avoid consequences of noncompliance.
@@
Although several licenses have been widely used for many years, interpretive questions still exist. Besides the preliminary questions about the types of use a license permits or prohibits, questions can also arise about whether two licenses are compatible with each other, as well as how certain terms apply to desired software uses. For example, the requirements of GPL v2 apply to any work “based on the Program,” which the license defines as equivalent to “any derivative work under copyright law.” But this language does not answer the question of whether, for example, system calls to a library create a work “based on the Program” or whether the answer depends on whether the link is static or dynamic, a question that even OSS lawyers have identified as unresolved  (Wintersgill et al., 2024).
2.2. Developers performing licensing tasks
@@
Previous work has shown that software developers often find licensing tasks difficult, such as determining whether two licenses are compatible with each other (Almeida et al., 2019, 2017), selecting or changing licenses for their software (Vendome et al., 2015), ensuring that their software complies with all relevant licensing requirements (Vendome et al., 2018), and releasing their software under multiple licenses (Moraes et al., 2021). Building on this work, we seek to contextualize these challenges by examining how often they are encountered and how developers address them.
@@
Both open source and proprietary tools can help with license compliance, such as Fossology (14) and Black Duck (5), respectively. Software bills of materials (SBOMs) can provide insights into software’s composition, facilitating licensing tasks (Bi et al., 2024; Stalnaker et al., 2024; Xia et al., 2023). Additionally, researchers have developed a variety of tools and processes aimed at helping developers with compliance tasks (Tuunanen et al., 2009; Ombredanne, 2020; Kapitsaki et al., 2017; German and Hassan, 2009; German and Di Penta, 2012; German et al., 2010; Hemel et al., 2011; Feng et al., 2019; Gangadharan et al., 2008). Some of these studies have explored the use of generative AI to assist in compliance tasks (Cui et al., 2025; Kahol et al., 2025; Ke et al., 2025), although the effectiveness of such tools still needs to be further evaluated. Other studies, based on mined software repositories and Q&A websites, have explored methods to detect non-approved licenses, non-standard license variants, and exceptions (Meloca et al., 2018; Zacchiroli, 2022; Papoutsoglou et al., 2022; Vendome et al., 2017b; Xu et al., 2023).
@@
Prior work has explored some of the decisions developers make during license compliance efforts. Wu et al. conducted a large empirical study aimed at understanding license usage (Wu et al., 2024). Works by Vendome et al. and Di Penta et al. sought to understand when, how, and why developers opt to change the license of their software (Di Penta et al., 2010; Vendome et al., 2017a). Liu et al. introduced a method to predict license changes from source code changes (Liu et al., 2019). Our previous work surveyed legal professionals with experience in software license compliance to understand their challenges and their views on software license compliance tasks (Wintersgill et al., 2024). Lastly, Li et al. provide a systematic literature review of stakeholders’ challenges in license compliance tasks (Li et al., 2025).
@@
Whereas the literature offers insights, solutions, and tools for licensing tasks, this work seeks to expand on previous efforts in two key ways: (1) by understanding if and how developers adopt these tools and techniques in practice, and (2) by identifying the license compliance pain points that are most relevant to developers by providing greater insight into the specific problems developers report encountering most often, the problems they report as being most difficult, and the strategies they use to address them.
@@ 3. Study Methodology
The goal of this study is to analyze the processes employed and challenges faced by software developers in performing OSS license compliance tasks to facilitate an understanding of the areas developers struggle with most and enable targeted work developing solutions to such pain points. The study was conducted through a survey and follow-up interviews and involved a collaboration between software engineering (SE) and law researchers. The context of the study consists of 58 developers recruited through personal contacts and open-source projects’ discussion channels.
The study aims to address the following research questions (RQs):
@@
    Which licensing tasks do developers struggle with? This RQ explores the licensing tasks developers encounter most frequently and which are the most difficult to complete.
@@
    What are the processes by which software licensing problems are resolved? This RQ explores how developers resolve compliance issues when they occur and investigates how developers perceive the relationship between development and legal teams.
@@
    How are developers incorporating license compliance tools in their workflows, if at all, and what issues have they encountered? This RQ explores how licensing tools are used by developers, if at all, and issues developers report when using such tools.
This study, including the survey questionnaire, the participant identification procedure, and the survey and interview protocols, was approved by our institution’s ethics review board. An overview of our methodology is depicted in Figure 1.
@@ 3.1. Survey design and participant identification
When designing our survey questionnaire, we considered previous literature on software license compliance, as well as general guidelines (Groves et al., 2009) and SE-specific guidelines for survey design (Pfleeger and Kitchenham, 2001; Kitchenham and Pfleeger, 2002a, c, b, d, 2003). The survey went through multiple review cycles, involving seven SE researchers and one law researcher, to achieve appropriate content, provide clear and concise language, and eliminate confusion and bias. We tried to limit the response time of the survey to 10 minutes or less to minimize the cognitive load on the participants and to encourage higher completion rates. In addition, we conducted a small pilot study with members of our research laboratory to further refine the survey questions, including by identifying questions that could be culled or reworked to better focus on the study’s research goals. The final survey included questions to obtain information related to each of our RQs, including questions relating to the compliance challenges faced by developers, existing compliance practices, and the use of tooling or other automation to perform compliance. Figure 2 provides an overview of our survey questionnaire. The complete survey text and information necessary for analysis can be found in our online replication package (Nathan Wintersgill, Trevor Stalnaker, Daniel Otten, Laura A. Heymann, Oscar Chaparro, Massimiliano Di Penta, Daniel M. German, Denys Poshyvanyk, 2026). (Compliance with our ethics board’s requirements prevents us from supplying all data collected.)
To recruit developers with diverse experiences, we distributed the survey (1) through our professional networks (posts on LinkedIn and emails), (2) through snowball sampling, encouraging respondents to share the survey within their networks, and (3) by posting links to the survey on relevant servers on Discord. We considered various other options for survey distribution (e.g., posting on relevant subreddits), but the selected methods allowed us to reach relevant communities while mitigating the risk of receiving responses from people outside the target population, which might occur more from survey invitations posted in highly public forums.
To identify Discord communities, we examined the TIOBE index (Jansen, 2025) of top programming languages to identify the top 20 most popular programming languages as of January 2025. We then searched for Discord servers associated with these languages (excluding Scratch, which is primarily a tool for beginners) by locating server links through the languages’ own documentation or websites, through third-party Discord server browsing tools DiscordMe (DiscordMe, ) and Disboard (Disboard, ), and through posts on subreddits related to the targeted languages. This process was supplemented by adding three generalized programming/development-focused servers and one server each for the popular React and Node.js frameworks. This yielded 27 Discord servers representing 16 of the top programming languages as well as general programming spaces and popular frameworks. Next, one researcher carefully analyzed the rules of each Discord server to determine whether it would permit a solicitation for participation in a survey, asking server administrators in cases where an answer was not immediately clear. In total, we were able to post survey invitations in 14 software development-focused Discord servers, with total membership exceeding 485,000 users at the time of distribution. Two additional posts were made in smaller servers but were later removed by moderators.
Figure 1. Research methodology (image credits at (Nathan Wintersgill, Trevor Stalnaker, Daniel Otten, Laura A. Heymann, Oscar Chaparro, Massimiliano Di Penta, Daniel M. German, Denys Poshyvanyk, 2026))
@@ 3.2. Survey response collection and analysis
Survey responses were collected using Qualtrics (47) from February 2025 to June 2025. Survey invitations were sent out iteratively during this same period. To monitor where responses originated from, we distributed survey URLs with identifiers unique to each source of participants. We obtained 58 complete survey responses. Of these, 38 came from our professional networks and 17 came from Discord, while 3 participants seemingly removed the tracking component from the survey URL. The largest population from a single Discord server was 6 participants.
Responses to the five open-ended questions were analyzed through a qualitative coding approach (Spencer, 2009). Three authors, all SE researchers, performed open-coding by independently assigning one or more codes to each response using a shared codebook (i.e., a set of spreadsheets). Each annotator independently coded all 58 responses, adding new codes to the codebook as necessary. Once the initial coding was completed, the annotators met to settle disagreements and consolidate the set of codes by discussing each response for which disagreements occurred and mutually deciding on a final set of codes to apply. Reasons for disagreement included different interpretations of potentially ambiguous responses, applying the same codes in different contexts, and using broader codes rather than more specific ones. All such disagreements were resolved among the annotators, resulting in a final set of codes, definitions, and assignments for each response. Our replication package (Nathan Wintersgill, Trevor Stalnaker, Daniel Otten, Laura A. Heymann, Oscar Chaparro, Massimiliano Di Penta, Daniel M. German, Denys Poshyvanyk, 2026) contains our final codes and definitions. We did not base our analysis on inter-rater agreements, as the nature of the study was such that codes were developed through an inductive process, and multiple codes could be assigned to each survey response. However, we carefully followed the best open-coding practices (Spencer, 2009), and we leveraged coders’ discussions to ensure the reliability of the results.
We corrected grammar and spelling errors in quotations presented in Section 4 for readability and indicated by ellipses or brackets any deletions or additions to text for clarity or space-related reasons. All responses have been attributed to survey participants using anonymous IDs for traceability (e.g., $R_{12}$).